# DRAGON: LLM-Driven Decomposition and Reconstruction Agents for Large-Scale Combinatorial Optimization

### v1

@@ Abstract.
Large Language Models (LLMs) have recently shown promise in addressing combinatorial optimization problems (COPs) through prompt-based strategies. However, their scalability and generalization remain limited, and their effectiveness diminishes as problem size increases, particularly in routing problems involving more than 30 nodes. We propose DRAGON, which stands for Decomposition and Reconstruction Agents Guided OptimizatioN, a novel framework that combines the strengths of metaheuristic design and LLM reasoning. Starting from an initial global solution, DRAGON autonomously identifies regions with high optimization potential and strategically decompose large-scale COPs into manageable subproblems. Each subproblem is then reformulated as a concise, localized optimization task and solved through targeted LLM prompting guided by accumulated experiences. Finally, the locally optimized solutions are systematically reintegrated into the original global context to yield a significantly improved overall outcome. By continuously interacting with the optimization environment and leveraging an adaptive experience memory, the agents iteratively learn from feedback, effectively coupling symbolic reasoning with heuristic search. Empirical results show that, unlike existing LLM-based solvers limited to small-scale instances, DRAGON consistently produces feasible solutions on TSPLIB, CVRPLIB, and Weibull-5k bin packing benchmarks, and achieves near-optimal results (0.16% gap) on knapsack problems with over 3M variables. This work shows the potential of feedback-driven language agents as a new paradigm for generalizable and interpretable large-scale optimization.
@@
Combinatorial Optimization, Large-scale, Metaheuristics
Zhuoyi Lin is the corresponding author.
@@ 1. Introduction
Figure 1. Illustration of DRAGON on a TSP instance, where metadata represent the coordinates of locations to visit, serving as the main environment for the agents. The decomposer identifies a suboptimal active segment {4,6,7,8} (grayed) from a global solution. The segment is locally refined and reintegrated by the reconstructor to improve the global tour.
Combinatorial optimization problems (COPs), such as the Traveling Salesman Problem (TSP), Vehicle Routing Problem (VRP), and Knapsack Problem (KP), are notoriously challenging due to their NP-hard nature Ausiello et al. (2012). These COPs are typically addressed using exact algorithms or heuristic methods, which often require substantial manual design and parameter tuning Rokbani et al. (2021); Cacchiani et al. (2022); Ma et al. (2023); Lin et al. (2024); Feng et al. (2025), thereby limiting their scalability and adaptability to varying problem instances and domains.
Large Language Models (LLMs) have led to a series of breakthroughs for various challenging problems in natural language processing and artificial intelligence, including language understanding, text generation, code synthesis, and reasoning tasks (Brown et al., 2020; Chowdhery et al., 2023; Achiam et al., 2024). Recent evidence reveals that LLMs have emerged as powerful tools capable of addressing a variety of COPs, leveraging their reasoning capabilities for direct solution generation Yang et al. (2024); Wang et al. (2024) or heuristic design Ye et al. (2025); Romera-Paredes et al. (2024); Liu et al. (2024b); Zhu et al. (2025). Their inherent strengths in abstraction, generalization, and semantic understanding allow them to generate promising solutions even without explicit algorithmic instructions. However, despite these advances, the current capabilities of LLMs for directly generating solutions remain largely confined to relatively small-scale instances, such as the TSP with fewer than 30 nodes Yang et al. (2024); Iklassov et al. (2024); Liu et al. (2024a). As problem size and complexity grow, LLM-based solutions often deteriorate due to inherent limitations, including restricted context length, reduced logical coherence, and difficulty in representing combinatorial structures Iklassov et al. (2024); Zhao et al. (2025). These limitations significantly impede the practical deployment of LLM-driven methods in realistic large-scale applications, such as logistics, transportation, and supply chain management, where problems commonly involve hundreds to thousands of nodes Bengio et al. (2021).
@@
Prompt-based LLM methods often fail on large-scale instances due to the absence of domain knowledge and the context length limitation, while code-generation approaches require extensive task-specific training to develop robust heuristics, leading to substantial computational costs and poor cross-domain transferability. Meanwhile, classical metaheuristics such as divide-and-conquer and large neighborhood search (LNS) Pisinger and Ropke (2018) exhibit strong scalability in large-scale COPs by iteratively decomposing and refining solutions, though they rely heavily on handcrafted heuristics and expert knowledge. To address the gap, we propose DRAGON, a framework that transforms complex large COP solving into decomposition and reconstruction tasks guided by LLM agents. DRAGON first strategically identifies regions with high potential of improvement, decomposing large-scale COPs into context-manageable subproblems, and locally optimizes each subregion through reconstruction.
@@ 2. Related Work
LLMs for Combinatorial Optimization. Recent studies have increasingly examined the capability of LLMs to solve COPs by leveraging prompt engineering techniques. Among these efforts, a prominent direction involves using LLMs directly as solution generators, where carefully designed prompts and enriched input information guide the models to produce viable solutions. Early research by Wang et al. (2024) provided evidence that LLMs could solve certain graph-based COPs, although this capability was strongly tied to the quality of the provided prompts. Subsequent work has sought to refine this process by incorporating contextual information. Examples include integrating graph structural and topological data Wasserkrug et al. (2024), existing heuristic solutions Iklassov et al. (2024), explicit relationships between solutions and objectives Yang et al. (2024), and visual representations of problems or solutions Huang et al. (2024); Elhenawy et al. (2024a). Another research pathway involves leveraging LLMs as components within structured algorithms or frameworks. For instance, Liu et al. (2024a) demonstrated the utility of LLMs in evolutionary computation by prompting them to perform essential algorithmic operations such as selection, crossover, and mutation. Similarly, Elhenawy et al. (2024b) deployed multiple LLMs as collaborative agents, each responsible for distinct roles such as initialization, critique, and evaluation within a predefined optimization pipeline. Despite these promising advancements, current capabilities for directly generating solutions via LLMs remain largely restricted to relatively small-scale problems, typically limited to instances such as TSP with fewer than 30 nodes. To circumvent this limitation, Iklassov et al. (2024) introduced a decomposition approach wherein an LLM assesses problem complexity autonomously. If deemed difficult, the model recursively subdivides the problem into simpler, manageable subproblems. Nonetheless, its scalability remains limited.
@@
Beyond direct solution generation, other studies have explored leveraging LLMs for heuristic design Romera-Paredes et al. (2024); Liu et al. (2024b); Ye et al. (2025) and mathematical modeling Xiao et al. (2023); AhmadiTeshnizi et al. (2024); Jiang et al. (2025b). The former aims to use LLMs to discover heuristics expressed in code that have potential to outperform those crafted by human experts, while the latter focuses on translating natural language problem descriptions into mathematical formulations compatible with traditional OR solvers. More recently, research has investigated training large language models for end-to-end combinatorial optimization, rather than relying solely on prompting paradigms Jiang et al. (2025a, 2024). These directions fall outside the scope of this work, and we refer interested readers to Da Ros et al. (2025) for a comprehensive survey.
@@
Large-Scale Combinatorial Optimization. Another research direction related to this work aiming to design metaheuristics for solving large-scale COPs, either with machine learning Li et al. (2021); Sun and Yang (2023) or through traditional methods Shaw (1998); Arnold et al. (2019). These approaches generally follow decomposition or divide-and-conquer principles. While they have shown promising results, their effectiveness often relies on problem-specific strategies (e.g., domain-specific decomposition rules), which limits their generality. In this work, we propose DRAGON, a general decomposition-reconstruction framework leveraging the semantic capabilities of LLMs. Unlike previous methods, DRAGON identifies suboptimal regions without explicit domain-specific rules and iteratively reconstructs globally consistent solutions. Our framework effectively addresses LLM context-length constraints, combining learned decomposition advantages with the generalization provided by LLM agents.
@@ 3. Methodology
Figure 2. Overview of the state passing among agents in DRAGON framework. With (a) given COP data input and current global solution, the DRAGON pipeline process alternates between two key stages: (b) A decomposition step uses an LLM to split a large-scale COP into manageable active and static segments. These are compressed into a smaller subproblem; (c) A reconstruction step solves the reduced problem with additional constraints to yield a refined local solution. This is then reintegrated into the global solution.
DRAGON is a two-stage framework that enables LLM agents to effectively solve large-scale COPs via state passing communication. As illustrated in Figure 2, DRAGON iteratively alternates between two core tasks: Decomposition and Reconstruction. Starting from an initial solution generated by any fast heuristic, DRAGON progressively refines it by identifying subregions with high improvement potential and solving localized subproblems under explicit constraint awareness. In each iteration, the current solution is decomposed into manageable subproblems, locally optimized through reconstruction agents, and then reassembled into a globally improved solution. This hierarchical process unifies local and global optimization, allowing DRAGON to scale from small instances to complex large-scale COPs. The example workflow illustrated in Figure 1 is summarized in Algorithm 1, where implicit communication between agents occurs via solution state passing, and the subsequent sections detail each component of the framework.
@@ 3.2. Decomposition for Large-Scale COPs
A central challenge in applying LLMs to large-scale COP decomposition lies in their limited capability to reason over complex global constraints. Although existing studies demonstrate the potential of LLMs in solving COPs Wilson et al. (2023); Yang et al. (2024), their capability to accurately detect and isolate sub-problems suitable for local refinement within large-scale COPs remains largely unexplored. To this end, we propose a decomposition agent ($\mathcal{D}$) which leverages LLMs to strategically partition COPs into smaller, manageable sub-problems for further refinement, rather than directly solving the entire COP. Specifically, the decomposer $\mathcal{D}$ functions as a high-level planner, analyzing metadata $M$ and the current solution $S_{i}$. It partitions the solution into active segments ($\mathbf{a}_{i}$), which have substantial improvement potential, and static segments ($\mathbf{s}_{i}$), which remain unchanged during the current iteration. Specifically, decomposition can be formulated as:
@@
where $M^{\prime}_{i}\subseteq M$ denotes the subset of original metadata relevant only to the active segments. For example, in the TSP scenario illustrated in Figure 1, the decomposer identifies an active segment $\mathbf{a}_{i}=\{4,6,7,8\}$ and a static segment $\mathbf{s}_{i}=[9,0,1,2,3,5]$. The Compress step significantly reduces the size of the local data $|M_{i}^{\prime}|<<|M|$, making it manageable to downstream agents. However, it will introduce extra specific constraints, which are a set of compressed representations of static segment $\mathbf{s}_{i}$ (e.g., enforce edge preservation like $(9,5)\in C_{i}$).
@@ 3.3. Reconstruction with Constraint Integration
Recent studies Yang et al. (2024); Liu et al. (2024a); Iklassov et al. (2024) highlight the effectiveness of LLMs in solving small-scale COPs, typically with fewer than 30 decision variables. Building on this capability, our reconstruction agent ($\mathcal{R}$) is designed to solve compressed COP instances iteratively, derived from the decomposition stage.
@@ 4.1.1. Problem Descriptions
  • Traveling Salesman Problem (TSP): A classical combinatorial routing problem that seeks the shortest possible tour visiting each city exactly once and returning to the starting point.
  • Capacitated Vehicle Routing Problem (CVRP): Extension of TSP with multiple vehicles and demand constraints. The goal is to minimize total travel distance while satisfying delivery demands.
  • Bin Packing Problem (BPP): Classical combinatorial optimization problem where items of various sizes must be packed into a minimal number of fixed-capacity bins without exceeding their limits.
  • Multiple Knapsack Problem (MKP): A generalization of the 0–1 knapsack problem involving multiple independent knapsacks. Unlike the multi-dimensional knapsack problem Cacchiani et al. (2022), where each item is constrained by several resource limits, the MKP aims to maximize the total value by optimally assigning items to knapsacks without exceeding their individual capacities.
@@ 4.1.2. Datasets
  • TSP We use 77 benchmark instances from TSPLIB Reinelt (1991) of the EUC_2D type, with sizes ranging from 50 to 20k nodes, covering a broad range of scales and spatial layouts. Oracle: optimal values or lower bounds reported in Reinelt (1991).
  • CVRP We select subsets of 19 instances from CVRPLIB, including both X-type Uchoa et al. (2017) (up to 1000 nodes) and larger XML-type instances Queiroga et al. (2022) distribution, which was synthesized at large scale (up to 5k nodes) by Zhou et al. (2023). Oracle: optimal values or lower bounds reported in Uchoa et al. (2017); Zhou et al. (2023).
  • BPP We adopt instances from FunSearch Romera-Paredes et al. (2024) and EoH Liu et al. (2024b) of all Weibull-5k, which is reflecting real-world scheduling and allocation scenarios. Oracle: L2 lower bounds determined by Romera-Paredes et al. (2024).
  • MKP Following Google OR-Tools guidelines Perron and Furnon (2025), we generate 10 synthetic instances by with knapsack capacities uniformly sampled from $U(100,500)$, using 10 to 100 knapsacks. Item values and weights drawn uniformly from $U(1,100)$. All instances will be released in supplementary JSON files. Oracle: optimal values or uppers bound searched by solver.
@@ 4.1.3. Implementation Details
We evaluate our framework using four representative LLMs accessed via API: two general-purpose models, OpenAI’s gpt-4o and gpt-4.1, and two reasoning-specialized models, OpenAI’s o3 and DeepSeek’s r1. Due to time and cost constraints, we limit our experiments to these models. Prompting strategies for each stage of DRAGON are illustrated in Figure 2, with complete prompt templates provided in the Appendix (Prompt designs). We adopt simple global solution initialization algorithms to each problem: 1) Random Insertion Azar (1994) for TSP, 2) Greedy Nearest Neighbor heuristic Mohammed et al. (2017) for CVRP, and 3) First-Fit Decreasing Baker (1985) for both BPP and MKP. For comparative evaluation, we implement OR-Tools Perron and Furnon (2025) solvers as: CVRP is solved via the routing model Furnon and Perron (2025) with Guided_Local_Search [^1] https://developers.google.com/optimization/routing/routing_options as the metaheuristic, while BPP uses the CP-SAT Perron and Didier (2025). All experiments run on an Ubuntu 20.04 server with an Intel Core i9-10900X (20 cores at 4.6 GHz) without GPU. For detailed prompt designs used in the decomposition and reconstruction processes, please refer to the Supplementary Materials.

### v2

# arXiv:2601.06502v2[cs.AI] 02 Mar 2026

@@ Abstract.
Large Language Models (LLMs) have recently shown promise in addressing combinatorial optimization problems (COPs) through prompt-based strategies. However, their scalability and generalization remain limited, and their effectiveness diminishes as problem size increases, particularly in routing problems involving more than 30 nodes. We propose DRAGON, which stands for Decomposition and Reconstruction Agents Guided OptimizatioN, a novel framework that combines the strengths of metaheuristic design and LLM reasoning. DRAGON autonomously identifies regions in initial solution with high optimization potential and strategically decompose large-scale COPs into manageable subproblems. Each subproblem is then reformulated as a concise, localized optimization task and solved through targeted LLM prompting guided by accumulated experiences. Finally, the locally optimized solutions are systematically reintegrated into the global context to yield a significantly improved outcome. By continuously interacting with the optimization environment and experience memory, the agents iteratively learn from feedback, effectively coupling symbolic reasoning with heuristic search. Empirical results show that, DRAGON consistently produces feasible solutions on TSPLIB, CVRPLIB, and Weibull-5k bin packing benchmarks, and achieves near-optimal results (0.16% gap) on knapsack problems with over 3M variables. This work shows the potential of feedback-driven language agents as a new paradigm for generalizable and interpretable large-scale optimization.
@@
Agentic AI; Combinatorial Optimization; Metaheuristics
doi:
SBAY6258
@@ 1. Introduction
Figure 1. Illustration of DRAGON on a TSP instance. The decomposer identifies a suboptimal segment {4,6,7,8} (grayed) from a global solution. The segment is locally refined and reintegrated by the reconstructor to improve the global tour.
Combinatorial optimization problems (COPs), such as the Traveling Salesman Problem (TSP), Vehicle Routing Problem (VRP), and Knapsack Problem (KP), are notoriously challenging due to the NP-hard nature ausiello2012complexity. COPs are typically addressed using exact algorithms or heuristic methods, which often require substantial manual design and parameter tuning rokbani2021bi; cacchianiKnapsackProblemsOverview2022; ma2023learning; lin2024cross; feng2025lifelong, thus limiting scalability and adaptability to varying problem instances and domains.
Large Language Models (LLMs) have led to a series of breakthroughs for various challenging problems, including language understanding, text generation, code synthesis, and reasoning tasks (brown2020language; chowdhery2023palm; achiam2023gpt). Recent evidence reveals that LLMs have emerged as powerful tools capable of addressing a variety of COPs, leveraging their reasoning capabilities for direct solution generation yang2024large; wang2024can or heuristic design ye2024reevo; romera2024mathematical; liu2024evolution; zhu2025bridging. Their inherent strengths in abstraction, generalization, and semantic understanding allow them to generate promising solutions even without explicit algorithmic instructions. However, the current capabilities of LLMs for directly generating solutions remain largely confined to relatively small-scale instances, such as TSP with fewer than 30 nodes yang2024large; iklassov2024self; liu2024large. As problem size and complexity increase, LLM-based solutions often deteriorate due to restricted context length, reduced logical coherence, and difficulty in representing combinatorial structures iklassov2024self; zhao2025can. These limitations significantly impede the practical deployment of LLM-driven methods in real large-scale applications, such as logistics, transportation, and supply chain management, where problems commonly involve hundreds to thousands of nodes bengio2021machine.
@@
Prompt-based LLM methods often fail on large-scale instances due to the absence of domain knowledge and the context length limitation, while code-generation approaches require extensive task-specific training to develop robust heuristics, leading to substantial computational costs and poor cross-domain transferability. Meanwhile, classical metaheuristics such as divide-and-conquer and large neighborhood search (LNS) pisinger2018large exhibit strong scalability in large-scale COPs by iteratively decomposing and refining solutions, though they rely heavily on handcrafted heuristics and expert knowledge. To address the gap, we propose DRAGON, a framework that transforms complex large COP solving into decomposition and reconstruction tasks guided by LLM agents. DRAGON first strategically identifies regions with high potential of improvement, decomposing large-scale COPs into context-manageable subproblems, and locally optimizes each subregion through reconstruction.
@@ 2. Related Work
LLMs for Combinatorial Optimization. Recent studies have increasingly examined the capability of LLMs to solve COPs by leveraging prompt engineering techniques. Among these efforts, a prominent direction involves using LLMs directly as solution generators, where carefully designed prompts and enriched input information guide the models to produce viable solutions. Early research by wang2024can provided evidence that LLMs could solve certain graph-based COPs, although this capability was strongly tied to the quality of the provided prompts. Subsequent work has sought to refine this process by incorporating contextual information. Examples include integrating graph structural and topological data wasserkrug2024large, existing heuristic solutions iklassov2024self, explicit relationships between solutions and objectives yang2024large, and visual representations of problems or solutions huang2024multimodal; elhenawy2024eyeballing. Another research pathway involves leveraging LLMs as components within structured algorithms or frameworks. For instance, liu2024large demonstrated the utility of LLMs in evolutionary computation by prompting them to perform essential algorithmic operations such as selection, crossover, and mutation. Similarly, elhenawy2024visual deployed multiple LLMs as collaborative agents, each responsible for distinct roles such as initialization, critique, and evaluation within a predefined optimization pipeline. Despite these promising advancements, current capabilities for directly generating solutions via LLMs remain largely restricted to relatively small-scale problems, typically limited to instances such as TSP with fewer than 30 nodes. To circumvent this limitation, iklassov2024self introduced a decomposition approach wherein an LLM assesses problem complexity autonomously. If deemed difficult, the model recursively subdivides the problem into simpler, manageable subproblems. Nonetheless, its scalability remains limited.
@@
Beyond direct solution generation, other studies have explored leveraging LLMs for heuristic design romera2024mathematical; liu2024evolution; ye2024reevo and mathematical modeling xiao2023chain; ahmaditeshnizioptimus; jiang2025droc. The former aims to use LLMs to discover heuristics expressed in code that have potential to outperform those crafted by human experts, while the latter focuses on translating natural language problem descriptions into mathematical formulations compatible with traditional OR solvers. More recently, research has investigated training large language models for end-to-end combinatorial optimization, rather than relying solely on prompting paradigms jianglarge; jiang2024bridge. These directions fall outside the scope of this work, and we refer interested readers to da2025large for a comprehensive survey.
@@
Large-Scale Combinatorial Optimization. Another research direction related to this work aiming to design metaheuristics for solving large-scale COPs, either with machine learning li2021learning; sun2023difusco or through traditional methods shaw1998using; arnold2019efficiently. These approaches generally follow decomposition or divide-and-conquer principles. While they have shown promising results, their effectiveness often relies on problem-specific strategies (e.g., domain-specific decomposition rules), which limits their generality. In this work, we propose DRAGON, a general decomposition-reconstruction framework leveraging the semantic capabilities of LLMs. Unlike previous methods, DRAGON identifies suboptimal regions without explicit domain-specific rules and iteratively reconstructs globally consistent solutions. Our framework effectively addresses LLM context-length constraints, combining learned decomposition advantages with the generalization provided by LLM agents.
@@ 3. Methodology
Figure 2. Overview of the state passing among agents in DRAGON framework. With (a) given COP data input and current global solution, the DRAGON pipeline process alternates between two key stages: (b) A decomposition step uses an LLM to split a large-scale COP into manageable active and static segments. These are compressed into a smaller subproblem; (c) A reconstruction step solves the reduced problem with additional constraints to yield a refined local solution.
DRAGON enables LLM agents to effectively solve large-scale COPs via state passing communication. As illustrated in Figure 2, DRAGON iteratively alternates between two core tasks: Decomposition and Reconstruction. Starting from fast heuristic initialization, DRAGON progressively refines it by identifying regions with high improvement potential and solving localized subproblems under explicit constraint awareness. In each iteration, the current solution is decomposed into manageable subproblems, locally optimized through reconstruction agents, and then reassembled into a globally improvement. This hierarchical process allows DRAGON to scale from small instances to complex large-scale COPs. Algorithm 1 summarized our workflow, where implicit communication between agents occurs via solution state passing.
@@ 3.2. Decomposition for Large-Scale COPs
A central challenge in applying LLMs to large-scale COP decomposition lies in their limited capability to reason over complex global constraints. Although existing studies demonstrate the potential of LLMs in solving COPs wilson2023; yang2024large, their capability to accurately detect and isolate sub-problems suitable for local refinement within large-scale COPs remains largely unexplored. To this end, we propose a decomposition agent ($\mathcal{D}$) which leverages LLMs to strategically partition COPs into smaller, manageable sub-problems for further refinement, rather than directly solving the entire COP. Specifically, the decomposer $\mathcal{D}$ functions as a high-level planner, analyzing metadata $M$ and the current solution $S_{i}$. It partitions the solution into active segments ($\mathbf{a}_{i}$), which have substantial improvement potential, and static segments ($\mathbf{s}_{i}$), which remain unchanged during the current iteration. Specifically, decomposition is formulated as:
@@
where $M^{\prime}_{i}\subseteq M$ denotes the subset of original metadata relevant only to the active segments. For example, in the TSP scenario illustrated in Figure 1, the decomposer identifies an active segment $\mathbf{a}_{i}=\{4,6,7,8\}$ and a static segment $\mathbf{s}_{i}=[9,0,1,2,3,5]$. The Compress step significantly reduces the size of the local data $|M_{i}^{\prime}|<<|M|$, making it manageable to downstream agents. However, it will introduce extra specific constraints, which are a set of compressed representations of static segment $\mathbf{s}_{i}$.
@@ 3.3. Reconstruction with Constraint Integration
Recent studies yang2024large; liu2024large; iklassov2024self highlight the effectiveness of LLMs in solving small-scale COPs, typically with fewer than 30 decision variables. Building on this capability, our reconstruction agent ($\mathcal{R}$) is designed to solve compressed COP instances iteratively, derived from the decomposition stage.
@@ 4.1.1. Problem Descriptions
$\bullet$ Traveling Salesman Problem (TSP): A classical combinatorial routing problem that seeks the shortest possible tour visiting each city exactly once and returning to the starting point.
$\bullet$ Capacitated Vehicle Routing Problem (CVRP): Extension of TSP with multiple vehicles and demand constraints. The goal is to minimize total travel distance while satisfying delivery demands.
$\bullet$ Bin Packing Problem (BPP): Classical combinatorial optimization problem where items of various sizes must be packed into a minimal number of fixed-capacity bins without exceeding their limits.
$\bullet$ Multiple Knapsack Problem (MKP): A generalization of the 0–1 knapsack problem involving multiple independent knapsacks. Unlike the multi-dimensional knapsack problem cacchianiKnapsackProblemsOverview2022, where each item is constrained by several resource limits, the MKP aims to maximize the total value by optimally assigning items to knapsacks without exceeding their individual capacities.
@@ 4.1.2. Datasets
Detailed datasets for the aforementioned problems:
$\bullet$ TSP We use 77 benchmark instances from TSPLIB reinelt1991tsplib of the EUC_2D type, with sizes ranging from 50 to 20k, covering a broad range of scales and spatial layouts. Oracle: optimal values or lower bounds reported in reinelt1991tsplib.
$\bullet$ CVRP We select subsets of 19 instances from CVRPLIB, including both X-type UCHOA2017845 (up to 1k nodes) and larger XML-type instances queiroga2022 distribution, which was synthesized at large scale (up to 5k nodes) by zhou2023towards. Oracle: optimal values or lower bounds reported in UCHOA2017845; zhou2023towards.
$\bullet$ BPP We adopt instances from FunSearch romera2024mathematical and EoH liu2024evolution of all Weibull-5k, which is reflecting real-world scheduling and allocation scenarios. Oracle: L2 lower bounds determined by romera2024mathematical.
$\bullet$ MKP Following Google OR-Tools guidelines ortools, we generate 10 synthetic instances by with knapsack capacities uniformly sampled from $U(100,500)$, using 10 to 100 knapsacks. Item values and weights drawn uniformly from $U(1,100)$. All instances will be released in supplementary JSON files. Oracle: optimal values or uppers bound searched by solver.
@@ 4.1.3. Implementation Details
We evaluate our framework using four representative LLMs accessed via API: two general-purpose models, OpenAI’s gpt-4o and gpt-4.1, and two reasoning-specialized models, OpenAI’s o3 and DeepSeek’s r1. Due to time and cost constraints, we limit our experiments to these models. Prompting strategies for each stage of DRAGON are illustrated in Figure 2, with complete prompt templates provided in the Appendix (Prompt designs). We adopt simple global solution initialization algorithms to each problem: 1) Random Insertion azar1994lower for TSP, 2) Greedy Nearest Neighbor heuristic mohammed2017solving for CVRP, and 3) First-Fit Decreasing baker1985new for both BPP and MKP. For comparative evaluation, we implement OR-Tools ortools solvers as: CVRP is solved via the routing model ortools_routing with Guided_Local_Search as the metaheuristic, while BPP uses the CP-SAT cpsatlp. All experiments run on an Ubuntu 20.04 server with an Intel Core i9-10900X (20 cores at 4.6 GHz) without GPU. For detailed prompt designs used in the decomposition and reconstruction processes, please refer to the Supplementary Materials.